Calculate your income elasticity of demand as your income increases from 10 000 to 12 000 if the price is 12, and the price is16. 


1
Expert's answer
2021-09-23T17:37:24-0400

Solution:

At the price of $12, when income increases from $10,000 to $12,000, the quantity demanded increases from 24 to 30 units.


Income elasticity of demand = "=\frac{\%\;change\; in\; quantity\; demanded}{\%\; change\; in\; Income}"


YEd = "\frac{Q_{2} -Q_{1}}{(Q_{2}+Q_{1})/2 } \div \frac{I_{2} -I_{1}}{(I_{2}+I_{1})/2 }"


Q1 = 24                                                I1 = 10,000

Q2 = 30                                                I2 = 12,000


YEd = "\frac{24 - 30}{(24+30)/2 } \div \frac{12,000 -10,000}{(680+440)/2 }"


= "\frac{6}{27 } \div \frac{2,000 }{11,000 } = \frac{0.22}{0.18 }= 1.22" 6/27 / 2,000/11,000 = 0.22/0.18 = 1.22


Income elasticity of demand (YEd) = 1.22

 

At the price of $16, when income increases from $10,000 to $12,000, the quantity demanded increases from 8 to 12 units.


Income elasticity of demand = "=\frac{\%\;change\; in\; quantity\; demanded}{\%\; change\; in\; Income}"


YEd = "\frac{Q_{2} -Q_{1}}{(Q_{2}+Q_{1})/2 } \div \frac{I_{2} -I_{1}}{(I_{2}+I_{1})/2 }"


Q1 = 8                                               I1 = 10,000

Q2 = 12                                             I2 = 12,000


YEd = "\frac{12 - 8}{(12+8)/2 } \div \frac{12,000 -10,000}{(12,000+440)/2 }"


= "\frac{4}{10 } \div \frac{2,000 }{11,000 } = \frac{0.4}{0.18 }= 2.22"


Income elasticity of demand (YEd) = 2.22